Please help me solve using the principle of zero products:
(9c+5)(4c-12)=0

Answer by mananth(16949)

(9c+5)(4c-12)=0
either (9c+5)=0
or (4c-12)=0
..
If (9c+5)=0
9c=-5
c=-5/9
If (4c-12)=0
4c=12
c=3